OpenStudy (jenny12):

A 14 foot tree casts an 8 foot shadow. At the same time, a nearby flagpole casts a 10 foot shadow. How tall is the flagpole?

OpenStudy (anonymous):

set up a proportion with the like units across from each other

OpenStudy (anonymous):

14 x --- + --- 8 10

OpenStudy (anonymous):

Assuming this is just a proportion problem, 14/8 = x/10. x=140/8

OpenStudy (jenny12):

Thanks.